OpenStudy (anonymous):

if f(x)=5x+4 and G(x)=6-3x calculate the following (f+g)(x)=

OpenStudy (anonymous):

\[(f+g)(x)=f(x)+g(x)=5x+4+6-3x=?\]does that help?

OpenStudy (anonymous):